A speaker generates sound waves which complete one vibration every 3 ms. Some person 170 metres away from the speaker hears the sound after 0.5 s from producing the sound. Calculate the distance between a compression and a successive rarefaction.


1
Expert's answer
2021-06-21T16:02:39-0400
"\lambda=\frac{v}{f}\\\lambda=\frac{L}{ft}
\\\lambda=\frac{170\cdot 0.003}{0.5 }=1\ m"